OpenStudy (taepoohwae):

expand the polynomial (7x+4)+(-8x+3)

OpenStudy (anonymous):

(7x+4)+(-8x+3)=7x+4-8x+3 (As +* - = - ie plus multiplies minus = minus) =7x-8x+4+3 =-x+7 is your required answer..